OpenStudy (anonymous):

Can anyone help me set up this problem? When a popular brand of cd player is priced at $300 per unit, a store sells 15 units per week. Each timethe price is reduced by $10, however, the sales increase by 2 per week. what selling price will result in weekly revenues of $7000

OpenStudy (anonymous):

So if we have $300/unit and are selling 15 units, then we are making a revenue of: $4500/wk. If we sell at $290/unit and are selling 17 units, we are making: $4930/wk. 250 * $25 = $6250/wk 230 * $29 = $6670/wk 200 * $35 = $7000/wk Hope this helps. Process of elimination. Facebook: Tutor Sean http://www.tutorsean.net

OpenStudy (anonymous):

thanks, i was hoping to find a way to set up a polynomial so i could just solve for the 7k mark. i just can't figure it out

OpenStudy (anonymous):

If you want, like my facebook page, and there is a place where you can talk to other professional tutors as well. There are 4 of us, and one of them (Robert) is great at that sort of thing. :)